Solve the system by addition.
5x – 3y = 13
4x – 3y = 11

Answer by tommyt3rd(5050)

5x – 3y = 13
-4x + 3y = -11
x=2

5(2)-3y=13
-3y=13-10
y=-1
Sol: (2,-1)